OpenStudy (anonymous):

Solve for values in the range of 0

OpenStudy (anonymous):

\[\cos x-sinx=0\]

OpenStudy (anonymous):

Would it be safe to assume it's when the values are equal?

OpenStudy (anonymous):

Like at pi/4, etc.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

don't forget \(\frac{5\pi}{4}\) where they are also equal

OpenStudy (anonymous):

Thanks. :)

OpenStudy (anonymous):

yw